From GlueXWiki
Jump to: navigation, search

January 3, 2013 FDC meeting


  1. Production Construction Tracking (Dave),picture of the three packages
  2. Corrosion tests (Lubomir)
  3. Engineering (Bill)
  4. Electronics (Chris, Nick)
  5. Tests at EEL126 FDC E-log
  6. Other


Participants: Eugene, Bill, Dave, Chris, Nick, and Lubomir.


- Picture done before Christmas of the the packages #4,1,3 (front to back) is linked above. Now we have also the colling loops installed on package #4, HV wiring done, tomorrow Chris and Nick will finish the HV connectors and the package will be ready for testing at EEL126. What's not in the picture: package #2 with three cells assembled. All four packages were flushed with Nitrogen over the holidays and yesterday we switched the gas to Ar/CO2 mixture for package #2, will start testing it next week.

- Dave: working on the last two cathodes (#75-76) needed for the last production package #2: installing, testing, gluing daughter cards on them. Also working on three spare cathodes. Casey started gluing a new spare wire frame; glued the PCB ring, now ready for lamination. We have PCBs for one more wire frame. Mike will no longer work for us (accepted a position at the cryo-group), still will have enough manpower (Dave) to finish the FDC production.

- We had issues with the HVAC for the clean room starting yesterday in the morning (most likely) and this morning a steam hose in the HVAC broke that was replaced by Dave. Ron Bartec said that due to power glitch the AC stopped working. Since the humidity in the bay area was low enough we never got humidity above 43% inside the clean room. The temperature there didn't exceed 74 deg. They fixed the problem and now everything is going back to normal.

Cathode corrosion test

- Lubomir: test #3 (Ar/CO2 gas, at ~70degC) is completed, see the final spreadsheet attached. Before Christmas the second EPDM sample failed, we stopped the test and opened the vacuum chamber. It turned out all the samples have low resistance at the place of the O-rings and all the samples that failed were due to higher resistance on the trace next to the soldering point. The main conclusion from this test is that there are problems doing conventional soldering on 2 microns copper traces and that these problems appear most likely before the corrosion we want to study. Therefore we need different experimental set-up if we want to continue with these tests. Interestingly, we never had soldering problems when using Apiezon, probably it leaked over and covered the soldering place thus preventing its damage.

- Measuring the resistance on the trace where the O-ring sample was placed turned out to be tricky. This is due to a non-conductive layer that has developed somehow on the top of the copper close to the O-ring, so one needs a pin that will go through that layer but will make contact to the thin copper trace. Using this measurement technique Lubomir re-checked the samples that failed before. The EPDM sample from the second test that failed after 12 days at 100 deg C and 100% humidity turned out to be really corrosion below the EPDM material. However, the two EPDM samples from the first test that failed after 8-10 days showed low resistance at the trace below the EPDM; this has to be confirmed with Vlad since there were mistakes with sample labeling. In any case using the results from the three tests and assuming the activating energy is the same for the different conditions, one can put a lower limit for the corrosion problem to appear when using EPDM+Viton: 2.6 (-1.1/+2.4) years.


- Bill: working to reduce the price of the package spacers, didn't gain a lot, will stick to the old design. Fluorinert is here, the chiller will come Jan. 17.

- Bill is designing a set-up to test the cooling. Will use a tube to make one loop with 22 pre-amps clamped to the tube, powered with LV, and somehow insulated using styrofoam (Bill) or just covering the pre-maps with bags (Eugene) to simulate the lack of air-flow. Nick: the test can be done in 117, he has LV for 22 channels there.

- Bill also wanted to test the tool for the manifold installation on the cooling loops in the packages; it is used to cut the tubes at the right place and then install the manifolds. He wanted to try it first on cooling loops not installed on the package, but Dave suggested that we can do it on the third package since there in any case we have to remove the cooling loops (the tube splittings there are not at the correct positions).

Tests at EEL126

- The two AC units in 126 are now operational set at 70deg. However, the remote control is at the facilities so we have to make request each time we want to change the setting. The problem is that the temperature in 126 is a result of both, the AC and the heater working at the same time. When it's colder outside there's more heating and if the thermostat is in the AC unit it will not sense the temperature close to the heat blower that is just above the testing set-up.

- We are planing to move the fourth package to 126 an Monday afternoon. Dave will coordinate this with the truck driver on Monday.